OpenStudy (anonymous):

Find the area of the sector with a central angle of 120° and a radius of 8 inches. Leave in terms of π.

OpenStudy (jdoe0001):

\(\bf \textit{Area of a Sector of a Circle}=\cfrac{\theta \pi r^2}{360}\) r = radius \(\theta\) = the angle